# TF-Attack: Transferable and Fast Adversarial Attacks on Large Language Models

### v1

@@ 3.4 Multi-Disturb & Dynamic-Disturb
Algorithm 1 TF-Attack Algorithm
1:  Initialization: agent $\pi_{\theta}$ with parameters $\theta$, $\beta$, sentences number $M$
2:  for $i\leftarrow 1$ to M do
3:     using ensemble-prompt on GPT-3.5 to get Important Level $L$
4:     while not receive termination signal do
5:        for $t\leftarrow 1$ to K do
6:           get words $s_{t}1\backsim s_{t}\beta$ from $level_{t}$
7:           compute $\pi_{\theta}((a^{f}_{t},a^{s}_{t})|s_{t})\backsim\pi_{\theta}(a_{t}^{f}|s_{t})$
8:           Using SimCSE to get possible substitutes $s_{t}1\backsim s_{t}\beta$
9:           compute result scores $S_{t}$
10:           update $t\leftarrow t+1$
11:        end for
12:     end while
13:     initialize $G(\tau)$ 0
14:     for $j\leftarrow T$ to 1  do
15:        $G(\tau)\leftarrow\gamma G(\tau)+r_{j}$
16:        accumulate $J_{j}(\theta)$
17:     end for
18:     update $\theta\leftarrow\theta+\alpha\nabla J(\theta)$
19:  end for
3.5 Attack Algorithm
The attacking process is shown in Algorithm 1. Since $a_{t}^{f}$ is chosen through a probability distribution, the method is encouraged to explore more possible paths of substitutions. The instant result $r_{t}$ is obtained from victim model after once substitution actions.
In order to make the comparison fairer, we set the following constraints for TF-Attack as well as all baselines: (1) Max modification rate: To better maintain semantic consistency, we only keep adversarial samples with less than 40% of the words to be perturbed. (2) Part-of-speech (POS): To generate grammatical and fluent sentences, we use NLTK tools [^1] https://www.nltk.org/ to filter candidates that have a different POS from the target word. This constraint is not employed on BERT-Attack. (3) Stop words preservation: the modification of stop words is disallowed and this constraint helps avoid grammatical errors. (4) Word embedding distance: For Textfooler, we only keep candidates with word embedding cosine similarity higher than 0.5 from synonyms dictionaries Mrkšić et al. (2016). For mask-fill methods, following BERT-Attack, we filter out antonyms Li et al. (2020) via the same synonym dictionaries for sentiment classification tasks and textual entailment tasks.
Once the attack finish condition is meet, the method will terminate this current step and output the answer to check whether it is succeeded. The expected return of substitutions is defined as follows:
$J(\theta)=\mathbb{E}[G(\tau)]$ (3)
Thus the result is calculated by ouput of model and can be expressed as follows:
$\nabla J(\theta)=\frac{1}{M}\sum^{M}_{m=1}\nabla\log\pi_{\theta}(\tau^{(m)})G(\tau^{(m)})$ (4)
where $[\tau^{(1)},\tau^{(2)},...,\tau^{(M)}]$ are $M$ samples of trajectories. The discount factor $\gamma$ enables both long-term and immediate effects to be taken into account and trajectories with shorter lengths are encouraged.
@@ 4.1 Experimental Setups
We further manually validate the quality of the adversaries from three challenging properties. 1) Human prediction consistency (Con): how often human judgment aligns with the true label; 2) Language fluency (Flu): measured on a scale of 1 to 5 for sentence coherence Gagnon-Marchand et al. (2019); and 3) Semantic similarity ($\text{Sim}_{\text{hum}}$): gauging consistency between input-adversary pairs, with 1 indicating agreement, 0.5 ambiguity, and 0 inconsistency.
@@ 4.5 Manual evaluation
At the beginning of manual evaluation, we provided some data to allow crowdsourcing workers to unify the evaluation standards. We also remove the data with large differences when calculating the average value to ensure the reliability and accuracy of the evaluation results. We first randomly select 100 samples from successful adversaries in IMDB and MR datasets and then ask ten crowd-workers to evaluate the quality of the original inputs and our generated adversaries. The results are shown in Table 9. For human prediction consistency, humans can accurately judge 93% of the original inputs on the IMDB dataset while maintaining an 87% accuracy rate with our generated adversarial examples. This suggests that TF-Attack can effectively mislead LLMs without altering human judgment. Regarding language fluency, the scores of our adversarial examples are comparable to the original inputs, with a minor score difference of no more than 0.3 across both datasets. Moreover, the semantic similarity scores between the original inputs and our generated adversarial examples stand at 0.93 for IMDB and 0.82 for MR, respectively. The result on NLI task is in Table 10. Overall, TF-Attack successfully preserves these three essential attributes.
@@ 5.2 Adversarial Training
We further investigate to improve the robustness of victim models via adversarial training using the generated adversarial samples. Specifically, we fine-tune the victim model with both original training datasets and our generated adversaries and evaluate it on the same test set. As shown in Table 14, compared to the results with the original training datasets, adversarial training with our generated adversaries can maintain close accuracy, while improving performance on attack success rates, modification rates, and semantic similarity. The victim models with adversarial training are more difficult to attack, which indicates that our generated adversaries have the potential to serve as supplementary corpora to enhance the robustness of victim models.

### v2

@@ 4.1 Experimental Setups
Following (Fang et al., 2023), We further manually validate the quality of the adversaries from three challenging properties. 1) Human prediction consistency (Con): how often human judgment aligns with the true label; 2) Language fluency (Flu): measured on a scale of 1 to 5 for sentence coherence Gagnon-Marchand et al. (2019); and 3) Semantic similarity ($\text{Sim}_{\text{hum}}$): gauging consistency between input-adversary pairs, with 1 indicating agreement, 0.5 ambiguity, and 0 inconsistency.
@@ 4.5 Manual evaluation
We follow Fang et al., 2023 to perform manual evaluation. At the beginning of manual evaluation, we provided some data to allow crowdsourcing workers to unify the evaluation standards. We also remove the data with large differences when calculating the average value to ensure the reliability and accuracy of the evaluation results. We first randomly select 100 samples from successful adversaries in IMDB and MR datasets and then ask ten crowd-workers to evaluate the quality of the original inputs and our generated adversaries. The results are shown in Table 9. For human prediction consistency, humans can accurately judge 93% of the original inputs on the IMDB dataset while maintaining an 87% accuracy rate with our generated adversarial examples. This suggests that TF-Attack can effectively mislead LLMs without altering human judgment. Regarding language fluency, the scores of our adversarial examples are comparable to the original inputs, with a minor score difference of no more than 0.3 across both datasets. Moreover, the semantic similarity scores between the original inputs and our generated adversarial examples stand at 0.93 for IMDB and 0.82 for MR, respectively. The result on NLI task is in Table 10. Overall, TF-Attack successfully preserves these three essential attributes.
@@ 5.2 Adversarial Training
Following Fang et al., 2023, we further investigate to improve the robustness of victim models via adversarial training using the generated adversarial samples. Specifically, we fine-tune the victim model with both original training datasets and our generated adversaries and evaluate it on the same test set. As shown in Table 14, compared to the results with the original training datasets, adversarial training with our generated adversaries can maintain close accuracy, while improving performance on attack success rates, modification rates, and semantic similarity. The victim models with adversarial training are more difficult to attack, which indicates that our generated adversaries have the potential to serve as supplementary corpora to enhance the robustness of victim models.